# Visual CoT: Unleashing Chain-of-Thought Reasoning in Multi-Modal Language Models

### v1

@@ Abstract
This paper presents Visual CoT, a novel pipeline that leverages the reasoning capabilities of multi-modal large language models (MLLMs) by incorporating visual Chain-of-Thought (CoT) reasoning. While MLLMs have shown promise in various visual tasks, they often lack interpretability and struggle with complex visual inputs. To address these challenges, we propose a multi-turn processing pipeline that dynamically focuses on visual inputs and provides interpretable thoughts. We collect and introduce the Visual CoT dataset comprising 373k question-answer pairs, annotated with intermediate bounding boxes highlighting key regions essential for answering the questions. Importantly, the introduced benchmark is capable of evaluating MLLMs in scenarios requiring specific local region identification. Extensive experiments demonstrate the effectiveness of our framework and shed light on better inference strategies. The Visual CoT dataset, benchmark, and pre-trained models are available here to foster further research in this direction.
Keywords:
Multi-Modal Language Models Chain-of-Thought
00footnotetext: Corresponding author.
@@ 1 Introduction
With the success of large language models (LLMs) like GPT-4 [1] and Gemini [52], researchers are actively exploring ways to enhance these models by incorporating visual understanding capabilities. This enthusiasm has led to the emergence of multi-modal large language models (MLLM), including LLaVA [32, 33], SPHINX [13, 30], and Qwen-VL [3]. Involving the extraction of visual tokens from input images, these MLLMs mostly follow a two-stage schedule: first the alignment of these tokens with linguistic modalities, and then the joint processing in LLMs. MLLMs have demonstrated viability in various scenarios, such as image captioning, visual question answering, and optical character recognition, owing to their ability to generate plausible outputs and leverage the extensive knowledge of LLMs.
@@
However, most popular MLLMs are primarily trained to respond to instructions based on visual inputs, employing a decoder-only autoregressive design as a single black box. While these models exhibit impressive generation capabilities, they suffer from inaccurate information [29] and even hallucinations [14]. Moreover, the black-box design hinders the interpretability of visual-language models. Additionally, the potential of multi-turn in-context capability and the advantages of Chain-of-Thought [59, 72, 66] for LLMs have not been extensively explored in MLLMs. Some recent works, such as multimodal-CoT [73] and [65, 64], have shown improvements by incorporating text-level chain-of-thought reasoning or in-context learning. However, it remains uncharted whether existing MLLMs can benefit from chain-of-thought reasoning in the visual understanding process, along with their interpretability remains largely unexplored.
@@
Furthermore, humans comprehend intricate visual information differently, often by focusing on specific regions or details within a given sample. For instance, when asked for a detailed regional description, we tend to scan the entire image first, locate the references, and then focus on the targets. In contrast, most MLLMs process aligned image contexts in a fixed-grain manner with a large amount of compute  (e.g., CLIP [46], EVA2-CLIP [51], InternVL [9]). To mimic human-like efficient reasoning behaviors, models need to identify regions containing essential visual details and dynamically zoom in for adjusted context, which current MLLMs struggle with, leading them to seek information primarily from the text modality.
@@
Therefore, there is a pressing need to develop methods that can handle dynamic, multi-turn, and focused visual inputs, while providing more interpretable stages of processing to enhance the efficacy and applicability of MLLMs. However, two significant challenges hinder the design of such pipelines: the lack of intermediate visual Chain-of-Thought supervision in existing visual question-answering (VQA) datasets, and the reliance of popular MLLM pipelines on static image context inputs.
@@
To address these challenges, we propose a novel pipeline that unleashes the reasoning capability of MLLMs, along with the corresponding Visual Chain-of-Thought training dataset. Our system is designed to identify and output key regions in an image that provides detailed information relevant to a given question. It integrates its understanding of both the original image and the detailed local image to generate the final answer. To facilitate this research direction, we develop and release a Visual Chain-of-Thought dataset by annotating each visual question-answer pair with a bounding box highlighting the key region essential for answering the question. Remarkably, the obtained model named VisCoT achieves improved performance even without an additional visual Chain-of-Thought reasoning stage, which shows the benefits of this modeling process. Hence, we provide the corresponding visual Chain-of-Thought benchmark and pre-trained models for reproducibility, with the hope of fostering further research in visual Chain-of-Thought for MLLMs.
@@
  • We present a Visual Chain-of-Thought dataset comprising 373k data items, each consisting of a question, an answer, and an intermediate bounding box as CoT contexts. The dataset spans across five distinct domains, ensuring a rich variety of visual data styles.
@@
  • We conduct extensive experiments to demonstrate the effectiveness of the proposed framework and analyze different components and strategies to shed light on ongoing research in this direction.
@@ 2.1 Multi-modal LLMs
Since the advent of large language models (LLMs), their success in various language applications has paved the way for the development of multi-modal large language models (MLLMs), which aim to integrate vision and language modalities. Initially, MLLMs were treated as dispatch schedulers to connect vision expert models, such as VisualChatGPT [60], HuggingGPT [48], LMDrive [47], and MM-REACT [65], in order to extend language models to other tasks and modalities. More recently, MLLMs have focused on aligning these two modalities by leveraging extensive training on image-caption pairs or image-question conversations. Notable methods like LLaVA [33] train a projector that maps image tokens to aligned representations of pre-trained LLMs. Other approaches, such as BLIP-2 [26, 25], adopt a query transformer (Q-Former) to learn image embeddings using learnable queries after obtaining image features. In terms of training strategy, recent works [33, 3, 57, 74, 7] commonly employ a 2-stage framework. The first stage involves pre-training on image-caption pairs, while the second stage focuses on alignment by using question-answering triplets. MLLMs have also been extended to various applications, including fine-grained localization [58, 23] such as object detection [69], video understanding [68, 28, 8], and image generation [19, 45].
2.2 Reasoning Capability of LLMs and MLLMs
LLMs have demonstrated impressive reasoning capabilities, enabled by in-context learning (ICL)[4], which allows feeding prompted samples and context. This capability has been further enhanced by Chain-of-Thought (CoT)[59] prompting, which enables LLMs to generate coherent intermediate reasoning steps toward the final answer. Previous studies have shown that LLMs benefit from manually written demonstrations [59] as well as zero-shot prompting outputs [20].
@@ 2.2 Reasoning Capability of LLMs and MLLMs
However, due to the domain gap between vision and text data, MLLMs fail to naturally inherit this reasoning capability. To address this limitation, researchers have focused on enhancing the reasoning capability of MLLMs in both the training and prompting paradigms. For instance, Flamingo [2] bridges the gap between these two modalities by pre-training on interleaved visual and textual data. Similarly, other works leverage visual grounded-reasoning data in training, such as Shikra [6] and KOSMOS-2 [42]. More recently, V[61] and CogCoM[44] modify the general mechanism in MLLMs and collect a series of visual reasoning steps as training data. On the other hand, studies have also explored prompting models [15, 70, 71] to understand complex visual scenes and tasks, focusing on the details of prompting techniques in MLLMs.
Figure 1: Examples of the collected data with corresponding question-answer annotations and visual CoT bboxes. The red bboxes in the images highlight the critical image regions that provide necessary and related information for answering the questions.
3 Method
In this section, we present the methodology of our visual Chain-of-Thought MLLM (VisCoT). We begin by describing the data production process in Sec. 3.1, which involves creating a diverse range of data samples, each consisting of a question, answer, and a corresponding visual bounding box, across various domains. This process involves linguistic and visual annotators who collaborate to create question-answer pairs and provide intermediate chain-of-thought bounding boxes, specifying the corresponding region in the image for answering the question. We then delve into the construction and evaluation of the CoT benchmark in Sec. 3.2, providing detailed discussions into its development. Subsequently, in Sec. 3.3 and Sec. 3.4, we outline our CoT pipeline and the corresponding model training procedure. Specifically, we employ a compatible approach to train a general two-turns MLLM with CoT, which means our model is capable of performing inference both with and without the visual CoT process, making it adaptable to a broad range of applications.
3.1 CoT Data Production
Figure 2: Density distribution of the visual CoT dataset in terms of the relative proportion of bbox region to full image size. Different colors denote different source datasets. We find that text-oriented source datasets typically show smaller ratios compared to others.
Our objective is to develop a two-round dialogue pipeline for multi-modal large language models (MLLMs) that enables them to identify specific regions in an image requiring additional attention for improved response performance. Our MLLM, named VisCoT, can incorporate both the overall image context and localized regions in its analysis. To equip the MLLM with CoT capabilities, we curate a visual CoT dataset, as outlined in Tab. 1, sourced from 10 existing datasets spanning five distinct domains. Fig. 1 showcases representative examples from this dataset, highlighting the diverse range of images included. In Fig. 2, the majority of the annotated key regions occupy only a small portion of the entire image, highlighting the importance of identifying these crucial areas for enhancing the accuracy of responses. For the linguistic annotation component, we employ GPT-4 [1], a large language model renowned for its robust language understanding and generation capabilities. In the subsequent sections, we will elaborate on the meticulous generation methods employed for each domain-specific dataset.
Text/Doc. We choose three text-related datasets to create data in this domain: TextVQA [50], TextCaps [49], DocVQA [40]. The three datasets focus on text recognition and comprehension in a variety of images and documents. TextVQA and DocVQA have already provided question-answer pairs, which we directly utilized. TextCaps, providing only captions and OCR tokens, required us to employ a linguistic annotator to create corresponding questions and answers, with further details available in the appendix. For the visual CoT bounding box, PaddleOCR is employed to detect the image and align the detected words and sentences with the answers, using the OCR-identified regions that match the answers as the CoT bounding boxes. This process ensures that the areas highlighted by the bounding boxes are directly relevant to the questions posed.
Fine-Grained understanding. For this domain, we use Birds-200-2011 [55] which is a widely-used dataset for fine-grained visual categorization. This dataset is not only rich in visual data but also includes detailed annotations about various bird parts and their attributes, along with bird bounding boxes in each picture. To leverage this dataset for our MLLM, we have formulated questions that challenge the model to identify specific characteristics or features present in the birds. These questions are designed to test the MLLM’s ability to discern and recognize fine-grained details in the images.
General VQA. In Flickr30k [43] dataset, each image encompassed five captions and the bounding boxes of most objects mentioned in the caption are also annotated. Employing a similar approach to TextCaps, we use GPT-4 to generate questions that require focusing on small objects in the images. The visual CoT bounding boxes in our proposed dataset correspond to the bboxes of objects identified and annotated in the official dataset.
Chart. We select the InfographicsVQA [39] dataset for its high-resolution infographics, which are advantageous for training MLLMs to pinpoint answer locations. Like in our Text/Doc data, we apply OCR techniques to identify regions containing the answers, using these identified areas as the CoT bounding boxes for more precise model training.
Relation Reasoning. We select the Visual Spatial Reasoning (VSR) [31], GQA [17], and Open Images [22] datasets to construct a dataset focused on relation-reasoning. These datasets are rich in spatial relational information among objects in images. For our Chain of Thought (CoT) bounding boxes, we use the bounding boxes surrounding the objects relevant to the query. For instance, if the question is “What is the material of the desk left to the woman?”, the bounding box of the desk to the woman’s left is designated as the visual CoT bounding box, providing more visual context for the MLLM’s reasoning process.
Domain Source Dataset Size Used GPT-4? Description
Text/Doc TextVQA [50] 16k No Picture with text
TextCaps [49] 32k Yes Picture with text
DocVQA [40] 33k No Doc Picture
Fine-Grained understanding Birds-200-2011 [55] 4k No Picture of birds
General VQA Flickr30k [43] 136k Yes Picture
Chart InfographicsVQA [39] 15k No Infographic
Relation Reasoning VSR [31] 3k No Picture
GQA [17] 88k No Picture
Open images [22] 43k No Picture
@@ 3.1 CoT Data Production
Table 1: The overview of the visual CoT dataset. The dataset spans across five distinct domains and simultaneously encompasses a variety of source datasets, ensuring a broad representation of different styles of visual data.
3.2 Visual CoT Benchmark
In this section, we provide an overview of our visual CoT benchmark, which primarily focuses on scenarios where the MLLM needs to concentrate on specific regions within a complete image. We utilize 10 source datasets, as shown in Fig. 1, and when an official training/evaluation split exists, we adopt it. In cases where such a split does not exist, we randomly divide the dataset. Additionally, we incorporate the test split of SROIE [16] and DUDE [54] to evaluate the model’s zero-shot visual CoT capabilities.
Following the methodology of previous MLLM studies [27, 37], we employ ChatGPT [41] and ask it to assign a numerical score between 0 and 1, where a higher score indicates better prediction accuracy. For detailed information on the prompt used for ChatGPT-based evaluation, please refer to the appendix.
Figure 3: VisCoT Pipeline: VisCoT first extracts visual tokens from an image and pinpoints the key region relevant to the question. Then, it processes the localized visual information. Finally, the MLLM integrates the information from the overall and localized images to construct a comprehensive and accurate answer.
3.3 Framework
The primary aim is to enhance the MLLM with visual CoT capabilities, as depicted in Fig. 3. We choose the pre-trained ViT-L/14 of CLIP [46] as the vision encoder and Vicuna-7/13B [10] as our LLM, which has better instruction following capabilities in language tasks compared to LLaMA [53]. Consider an input original image $X_{0}$, we take the vision encoder to obtain the visual feature $Z_{0}$. Similar to LLaVA [33, 32], we use a simple linear layer to project the image features into the word embedding space to obtain the visual tokens $H_{0}$ which share the same dimensionality of the large language model.

### v2

# Visual CoT: Advancing Multi-Modal Language Models with a Comprehensive Dataset and Benchmark for Chain-of-Thought Reasoning

@@ Abstract
Multi-Modal Large Language Models (MLLMs) have demonstrated impressive performance in various VQA tasks. However, they often lack interpretability and struggle with complex visual inputs, especially when the resolution of the input image is high or when the interested region that could provide key information for answering the question is small. To address these challenges, we collect and introduce the large-scale Visual CoT dataset comprising 438k question-answer pairs, annotated with intermediate bounding boxes highlighting key regions essential for answering the questions. Additionally, about 98k pairs of them are annotated with detailed reasoning steps. Importantly, we propose a multi-turn processing pipeline that dynamically focuses on visual inputs and provides interpretable thoughts. We also introduce the related benchmark to evaluate the MLLMs in scenarios requiring specific local region identification. Extensive experiments demonstrate the effectiveness of our framework and shed light on better inference strategies. The Visual CoT dataset, benchmark, and pre-trained models are released to foster further research in this direction.
@@ 1 Introduction
With the success of large language models (LLMs) like GPT-4 [1] and Gemini [55], researchers are enhancing these models by incorporating visual understanding capabilities. This enthusiasm has led to the emergence of multi-modal large language models (MLLM), such as LLaVA [33, 34], SPHINX [14, 31], and Qwen-VL [3]. Involving the extraction of visual tokens from input images, these MLLMs mostly follow a two-stage schedule: first the alignment of these tokens with linguistic modalities, and then the joint processing in LLMs. MLLMs have demonstrated viability in various scenarios, such as image captioning, visual question answering, and optical character recognition, owing to their ability to generate plausible outputs and leverage the extensive knowledge of LLMs.
@@
However, many popular MLLMs are primarily trained to respond to instructions based on visual inputs, employing a decoder-only autoregressive design as a single black box. While these models exhibit impressive generation capabilities, they suffer from inaccurate information [30] and even hallucinations [15]. Moreover, the black-box design hinders the interpretability of visual-language models. Additionally, the potential of multi-turn in-context capability and the advantages of chain-of-thought [62, 75, 69] for LLMs have not been extensively explored in MLLMs. Some recent works, such as multimodal-CoT [76] and [68, 67], have shown improvements by incorporating text-level chain-of-thought reasoning or in-context learning. However, it remains uncharted whether existing MLLMs can benefit from chain-of-thought reasoning in the visual understanding process, along with their interpretability remains largely unexplored.
@@
Furthermore, humans comprehend intricate visual information differently, often by focusing on specific image regions or details within a given sample. For instance, when asked for a detailed regional description, humans tend to scan the entire image first, locate the references, and then focus on the targets. In contrast, most MLLMs process aligned image contexts in a fixed-grain manner with a large amount of computation  (e.g., CLIP [50], EVA2-CLIP [54], InternVL [9]). To mimic human-like efficient reasoning behaviors, models need to identify image regions containing essential visual details and dynamically zoom in to capture adjusted context, which current MLLMs struggle with, leading them to seek information primarily from the text domain.
@@
Therefore, there is a pressing need to develop methods that can handle multi-turn, dynamic focused visual inputs, while providing more interpretable stages of reasoning to enhance the efficacy and applicability of MLLMs. However, two significant challenges hinder the design of such pipelines: the lack of intermediate visual chain-of-thought supervision in existing visual question-answering (VQA) datasets, and the reliance of popular MLLM pipelines on static image context inputs.
@@
To address these challenges, we develop and release a 438k visual chain-of-thought dataset by annotating each visual question-answer pair with a bounding box. The bounding box highlights the key image region essential for answering the question. We suppose that accurately locating and comprehending this key region will significantly improve MLLM’s response accuracy and relevance. Notably, about 98k question-answer pairs include extra detailed reasoning steps. These annotations are designed to instruct the MLLM in a logical, step-by-step process to identify the final bbox and generate the answer. Building on the dataset, we propose a novel pipeline that unleashes the visual CoT reasoning capability of MLLMs, which is designed to identify and output key regions in an image that provides detailed information relevant to the given question. It integrates the understanding of both the original image and detailed local image to generate the final answer. Besides, we provide the corresponding visual CoT benchmark and pre-trained models for reproducibility, aiming to foster further research in the visual chain-of-thought for MLLMs.
Figure 1: Examples of five domains covered in the visual CoT dataset, with corresponding question-answer annotations and visual CoT bboxes: chart, text/doc, general VQA, fine-grained understanding, and relation reasoning. The red bounding boxes in the images highlight the critical image regions that provide necessary and related information for answering the questions.
@@
  • We present a visual chain-of-thought dataset comprising 438k data items, each consisting of a question, an answer, and an intermediate bounding box as CoT contexts. Some items also contain detailed reasoning steps. The dataset spans across five distinct domains.
@@ 2 Related Works
Multi-modal LLMs. Since the advent of large language models (LLMs), their success in various language applications has paved the way for the development of multi-modal large language models (MLLMs), which integrate vision and language modalities. Initially, MLLMs were treated as dispatch schedulers to connect vision expert models, such as VisualChatGPT [63], HuggingGPT [51], and MM-REACT [68], in order to extend language models to other tasks and modalities. More recently, MLLMs have focused on aligning these modalities through extensive training on image-caption pairs or image-question conversations. Notable methods like LLaVA [34] train a projector that maps image tokens to aligned representations of pre-trained LLMs. Other approaches, such as BLIP-2 [27, 26], adopt a query transformer (Q-Former) to learn image embeddings using learnable queries after obtaining image features. MoVA [82] designs an adaptive router to fuse task-specific vision experts with a coarse-to-fine mechanism. In terms of training strategy, recent works [34, 3, 60, 80, 7, 38] commonly employ a 2-stage framework. The first stage involves pre-training on image-caption pairs, while the second stage focuses on alignment by using question-answering triplets. MLLMs have also been extended to various applications, including fine-grained localization [61, 24] such as object detection [72], video understanding [71, 29, 8], and image generation [20, 49].
@@
Reasoning Capability of LLMs and MLLMs. LLMs have demonstrated impressive reasoning capabilities, enabled by in-context learning (ICL)[4], which allows feeding prompted samples and context. This capability has been further enhanced by chain-of-thought (CoT)[62] prompting, which enables LLMs to generate coherent intermediate reasoning steps toward the final answer. Previous studies have shown that LLMs benefit from manually written demonstrations [62] as well as zero-shot prompting outputs [21]. Trar [78] proposes a routing module to dynamically select informative regions based on the attention map. However, due to the domain gap between vision and text data, MLLMs fail to naturally inherit this reasoning capability. To address this limitation, researchers have focused on enhancing the reasoning capability of MLLMs in both the training and prompting paradigms. For instance, Flamingo [2] bridges the gap between these two modalities by pre-training on interleaved visual and textual data. Similarly, other works leverage visual grounded-reasoning [39, 79] data in training, such as Shikra [6] and KOSMOS-2 [46]. More recently, V[64] and CogCoM[48] modify the general mechanism in MLLMs and collect a series of visual reasoning steps as training data. On the other hand, studies have also explored prompting models [16, 73, 74, 44, 77] to understand complex visual scenes and tasks, focusing on the details of prompting techniques in MLLMs.
3 Visual CoT Dataset
There is a shortage of multimodal datasets for training multi-modal large language models (MLLMs) that require to identify specific regions in an image for additional attention to improve response performance. This type of dataset with grounding bbox annotations could possibly help the MLLM output intermediate interpretable attention area and enhance performance. To fill the gap, we curate a visual CoT dataset, as illustrated in Fig. 1 and Tab. 1. This dataset specifically focuses on identifying critical regions within images, a feature essential for models to concentrate on relevant visual elements to improve response accuracy. Each data sample consists of a question, answer, and a corresponding visual bounding box across five domains, as shown in Tab. 2. Some data samples also include extra detailed reasoning steps.
To ensure a robust foundation for detailed visual and textual analysis, our dataset deliberately integrates a diverse selection of data including text/doc, fine-grained understanding, charts, general VQA, and relation reasoning. These data domains are deliberately chosen to cultivate a comprehensive skill set across varied analytical tasks: 1) Text/doc enhances MLLM’s capabilities on OCR and contextual understanding, crucial for applications requiring text interpretation in complex environments. 2) Fine-grained understanding aids in identifying and distinguishing subtle differences in visual appearance and patterns. 3) Charts foster the ability to interpret graphical data, which are essential for business and scientific applications. 4) General VQA exposes models to a wide array of visual queries, improving their general usability. 5) Relation reasoning data develops spatial and contextual awareness of MLLMs, vital for interactive and navigational tasks. Together, these modalities ensure the dataset not only fills existing gaps but also enhances the versatility and contextual awareness of MLLMs across varied scenarios.
Table 1: One data example with detailed reasoning steps, of which we have collected about 98k of this type. The red bounding box shows the important image region for answering the question.
An example of detailed reasoning steps in GQA dataset
Question: What appliance is to the right of the cabinet? [Uncaptioned image]
Please think step by step and provide the bounding box coordinate
of the region that can help you answer the question better.
Reasoning steps: 1. Identify the cabinet in the image.
2. Observe the area to the right of the identified cabinet.
3. Look for any appliance located to the right side of the cabinet.
4. Determine the name of the appliance found in this location
CoT BBox: [163, 44, 206, 67]
Answer
The appliance is a microwave.
3.1 Data Generation
@@
Table 2: The overview of the visual CoT dataset. The dataset spans five distinct domains and includes various source datasets, ensuring a broad representation of visual data styles.
Domain Source Dataset Size Used GPT-4? Dataset Description
Text/Doc TextVQA [53] 16k No Images with text
TextCaps [52] 32k Yes Images with text
DocVQA [43] 33k No Doc Images
DUDE [57] 15k No Doc Images
SROIE [17] 4k No Invoice Images
Fine-Grained Understanding Birds-200-2011 [58] 10k No Images of birds
General VQA Flickr30k [47] 136k Yes Images
Visual7W [81] 43k No Images
Charts InfographicsVQA [42] 15k No Infographic
Relation Reasoning VSR [32] 3k No Images
GQA [18] 88k Yes Images
(with detailed reasoning steps)
Open images [23] 43k No Images
To collect and build a diverse and comprehensive Visual CoT dataset, we select twelve source datasets across five distinct domains, primarily consisting of Visual Question Answering (VQA) and Image Captioning datasets. We reuse their images and useful annotations, such as question-answer pairs, image captions, and object relations, to aid in building our dataset. The data construction process involves both linguistic and visual annotators to create question-answer pairs, and provide intermediate chain-of-thought bounding boxes indicating the crucial image region for answering the question. For the linguistic annotations, we employ GPT-4 [1], known for its robust language understanding and generation capabilities. For the visual annotations, we choose PaddleOCR [12], an efficient and accurate tool for optical character recognition. In the following sections, we elaborate on the generation methods employed for each domain-specific dataset.
Figure 2: Statistics of the proposed visual CoT dataset. We visualize the CoT bbox distribution, average bbox size, and average relative size of bbox area $R$ for each source dataset.
Text/Doc. We choose five text-related datasets to create data in this domain: TextVQA [53], DocVQA [43], DUDE [57], TextCaps [52], SROIE [17]. The five datasets focus on text recognition and comprehension in a variety of images and documents. TextVQA, DocVQA, DUDE and SROIE have already provided question-answer pairs, which we directly adopt. TextCaps, providing only captions and OCR tokens, required us to employ a linguistic annotator to create corresponding questions and answers (see further details in Appendix E.1.) For the visual CoT bboxes, we then apply PaddleOCR[12] to detect OCR-identified regions in the image, and specify the CoT bounding boxes as the region that consists of words and sentences aligning with the answer. Furthermore, we also design a filtering pipeline to improve content quality. This process ensures that the areas highlighted by the bounding boxes are directly relevant to the questions.
Fine-Grained Understanding. For this domain, we use Birds-200-2011 [58], which is a widely-used dataset for fine-grained visual categorization. This dataset is not only rich in visual data but also includes detailed annotations about various bird parts and their attributes, along with bird bounding boxes in each picture. To leverage this dataset for our MLLM, we have formulated questions that challenge the model to identify specific characteristics or features present in the birds. These questions are designed to test the MLLM’s ability to discern and recognize fine-grained details in the images.
General VQA. We use Flickr30k [47] and Visual7W [81] as the dataset for general VQA tasks. In Flickr30k, each image encompassed five captions and the bounding boxes of most objects mentioned in the captions. Employing a similar approach to TextCaps, we use GPT-4 to generate questions that require focusing on small objects in the images. The visual CoT bounding boxes in our proposed dataset correspond to the bboxes of objects identified and annotated in the official dataset. Visual7W has already provided the question-answer pairs with object-level grounding annotations.
Charts. We select the InfographicsVQA [42] dataset for its high-resolution infographics, which are advantageous for training MLLMs to pinpoint answer locations. Like in our Text/Doc data, we apply OCR techniques to identify regions containing the answers, using these identified areas as the CoT bounding boxes for more precise model training.
Relation Reasoning. We select the Visual Spatial Reasoning (VSR) [32], GQA [18], and Open Images [23] datasets to construct data focusing on relation-reasoning. These datasets are rich in spatial relational information among objects in images. For our chain-of-thought (CoT) bounding boxes, we use the bounding boxes surrounding the objects relevant to the question. For instance, if the question is “What is the material of the desk left to the woman?”, the bounding box of the desk to the woman’s left is designated as the visual CoT bounding box, providing more visual context for the MLLM’s reasoning process. In GQA [18] each image is associated with a scene graph of objects and relations. Each question comes with a structured representation of its semantics. With these annotations, we utilize GPT-4 to generate detailed reasoning steps, as illustrated in Tab. 1. The related prompt is available in Appendix E.3.
3.2 Dataset Analysis
We provide a visualization of the data statistics in Fig. 2. We partition the bboxes in each dataset into three groups (large, medium, small) based on the relative bounding box size $R$, which is the ratio of the CoT bbox size relative to the total image size. The visualization reveals that the majority of the annotated key regions, particularly in text-oriented datasets, occupy only a small portion of the entire image, highlighting the importance of identifying these crucial areas to enhance performance. Specifically, the average bounding box size is 247.82 pixels, which well aligns with the common input resolution for a vision encoder ranges between 224 and 336 pixels, while the original image size is usually too large and needs down-sampling that loses information. These regions account for only about 13.2% of the image area. This highlights the necessity for MLLMs to accurately pinpoint these crucial areas to enhance processing efficiency and effectiveness. If the model fails to correctly identify and focus on these key regions, the majority of the image processed could be irrelevant, leading to inefficient computation, hallucination, and potential degradation in performance.
4 Enhancing MLLMs with Chain-of-Thought Capabilities
Along with the visual CoT dataset, we also propose a visual CoT MLLM framework named VisCoT, which employs standard models without specialized modifications, serving as a baseline to enhance MLLMs with visual CoT capabilities. In this section, we briefly introduce the framework, and illustrate the pipeline in Fig. 3. Readers are referred to Appendix B for more details.